A student was told to make a pH = 7.5 buffer from sodium dihydrogen phosphate and sodium hydrogen phosphate with an ionic strength of 0.10

a) What concentration of Na2HPO4 and NaH2PO4 need to be used?

b) If 10.0 mL of 1.00 M Hbr is added to 1.00 L of this buffer, what is the pH?

Please show work.
Answer #1

TV o al. 4o 7.5 weak Acida na hapo Salt - Naz Poq Inic strength of salt = 0.1 Isot No, upog anat + ulog I= Ź [mxcht macete Is7:5_-7.31 Eloy [Salt) e cara Isait] 750-72 tua dag (soit E Aca] [sact] At Co.za) - S: 2 1.95 1.95 = Sait] Cara 1.95= 0.04 [Anie. Now New Conch. One Inateloi]e 0:02+10x1 003_mala [Noh pos] = 0.04-10x1 loo - 003: Now the New bh bu= bkat log [salt? Ava)